how many unis of computers o e gai how many units of cars do we lose? 00 7001,000Quantity of Cars Produced 300 2.2 How many cars are to be sacrificed to get one more unit of computer? which is asking the opportunity cost of computer at the point B in units of car What is the opportunity cost of computer at the point B in units of ear? cars 1-3. The farmer and rancher each work 8 hours/ day and can devote this time to growing potatoes, raising cattle, or a combination The farmer can produce an ounce of potatoes in 30 minutes and an ounce of meat in 60 minutes. The rancher can produce an ounce of potatoes in 10 minutes and an ounce of meat in 40 minutes Amounts produced in 8 hours (minutes for I oz.) potatoes meat Farmer Rancher 12 oz. (40min/ounce 48 oz (1Omin/ounce Opportunity Cost of .1. Answer the blanks in the table aside on the red-colored cells 1 oz. of meat oz. meat Farmer I)oz. meat 0.25 oz. meat much oz. potatoes 3.1.2 oz. potato Rancher (3.1.2 )oz. potatoes Chech up the answer 312 ght the nwern 1 the farmer (rancher) give up to produce in onder to produce 1 o of potatoesea 3.2 Who has an absolute advantage in producing potatoes Farmer? or Rancher? Answer on the red-colored cell. Note: compare amounts produced in a given time .3 Who has an comparative advantage in producing potatoes? Farmer? or Rancher? Answer on the red-colored cell. Note: compare efficieny (low costy in opportunity cost
Answer #1

Answer :

3.1.1

To produce 1 oz of potato farmer need to give up meat =8/16=0.5 oz

3.1.2

To produce 1 oz of meat rancher need to give up potato =48/16= 4 oz

3.2

Since rancher produces 48 oz of potato in 8 hrs compare to Farmer who only produces 16 oz so rancher has absolute advantage.

3.3

Since opportunity cost for producing 1 oz of potato for farmer is 0.5 where as for Rancher opportunity cost is 0.25. Since rancher has less opportunity cost.So Rancher has competitive advantage for producing potatoes.